# W4A16 Mixed-Precision Matrix Multiplication on Decoupled Architecture: Kernel Design and Memory Bottleneck Analysis for Ascend NPUs

### v1

@@ Abstract
As Large Language Models (LLMs) scale, weight-only quantization (W4A16: 4-bit weights, 16-bit activations) becomes critical for reducing memory footprint with minimal accuracy loss. However, its efficient deployment on Huawei’s Ascend 910 Neural Processing Unit (NPU) is challenging due to limited native mixed-precision support and the accelerator’s decoupled compute architecture. To enable quantization on such architecture, we present the first practical W4A16 matrix multiplication kernel tailored for the Ascend 910 NPU. Our design leverages vector cores for on-the-fly INT4-to-FP16 dequantization, cube cores for high-throughput GEMM, and Split-K parallelization to mitigate memory latency. Performance evaluations across diverse matrix shapes and batch sizes show our method outperforms data-parallel approaches when K N, a typical scenario in LLM decoding. Specially, our method can achieve a speedup ranging from to . In addition, our profile reveals the primary bottleneck is not dequantization compution itself, but extra global memory transfer for the weight, making W4A16 only reaching a maximum speedup of over native FP16FP16 matrix multiplication in PyTorch. In the long run, our method lays a solid foundation and provides insightful views for the efficient deployment of quantized large language models on various domain-specific accelerators.
@@ 1 Introduction
Quantization [7] has emerged as a pivotal technique for deploying large language models (LLMs) [13, 15, 9] on resource-constrained hardware. Recent studies, such as SINQ [10] and QQQ [16], have achieved significant reductions in memory footprint, bandwidth requirements, and computational cost without incurring substantial losses in model accuracy. Among various quantization schemes, weight-only strategies—wherein weights are quantized to fixed-point integers while activations remain in floating-point format—have garnered growing attention due to their capability to balance efficiency and fidelity. In particular, the W4A16 configuration (4-bit weights and 16-bit activations) [2, 8] has demonstrated remarkable potential: it retains high activation precision to preserve the dynamic range, while aggressively compressing weights to minimize storage and memory access overhead. However, unlocking the full potential of W4A16 quantization necessitates specialized kernel implementations that can efficiently unpack and dequantize 4-bit weights into FP16/BF16 formats prior to matrix multiplication. General Matrix Multiplication (GEMM) routines are ill-suited for such preprocessing-intensive, irregular data layouts and often fail to fuse dequantization with computation. This deficiency results in suboptimal memory bandwidth utilization and redundant data movement—challenges that are particularly pronounced on domain-specific architectures such as Huawei’s Ascend NPU [17]. While the cube units of the Ascend NPU are highly optimized for native GEMM operations, they lack tailored support for the additional memory-bound dequantization steps required by 4-bit weight formats.
@@ 2.2 Motivation and Challenges
W4A16[2, 8] quantization (4-bit weights, 16-bit activations) offers an trade-off between model efficiency and accuracy. By storing weights in 4-bit integers, it reduces model size by approximately 4× compared to FP16 representations. However, a key implementation challenge remains: mainstream hardware—including GPUs and NPUs—lacks native support for mixed-precision arithmetic operations. In practice, 4-bit values must be packed into larger data types such as INT32 for storage and transfer, then dequant prior to computation. This dequant process introduces additional overhead and can become a notable performance bottleneck unless carefully optimized through efficient pipline and memory access patterns. Thus, while W4A16 is promising for efficient inference, its practical speedup depends heavily on low-level kernel optimization and hardware-aware implementation.
@@ 4.1 Kernel Strategy Benchmarking
We evaluate the performance of INT4-weight × FP16-activation matrix multiplication kernel on the Ascend 910 NPU across a range of model dimensions and batch sizes, comparing two parallelization strategies: Split-K and Data Parallel. As shown in Figure 2, when is significantly larger than , the Split-K strategy outperforms data-parallel approaches, and our method achieves a speedup ranging from to . This is because Split-K can more effectively partition the computational workload across each cube core under such dimensional constraints. As the batch size (M) gradually increases, performance remains relatively stable. This occurs because the cube cores operate on matrix tiles as their fundamental processing granularity. Even when batch size is small, the input data is padded accordingly, which explains the limited acceleration benefit for small batch sizes. This behavior is observed in both Split-K and data-parallel methods.

### v2

@@ Abstract
As Large Language Models (LLMs) scale, weight-only quantization (W4A16: 4-bit weights, 16-bit activations) becomes critical for reducing memory footprint with minimal accuracy loss. However, its efficient deployment on Huawei’s Ascend 910 Neural Processing Unit (NPU) is challenging due to limited native mixed-precision support and the accelerator’s decoupled compute architecture. To enable quantization on such architecture, we present the first practical W4A16 matrix multiplication kernel tailored for the Ascend 910 NPU. Our design leverages vector cores for on-the-fly INT4-to-FP16 dequantization, cube cores for high-throughput GEMM, and Split-K parallelization to mitigate memory latency. Performance evaluations across diverse matrix shapes and batch sizes show our method outperforms data-parallel approaches when K N, a typical scenario in LLM decoding. Specially, our method can achieve a speedup ranging from to . In addition, our profile reveals the primary bottleneck is not dequantization computation itself, but extra global memory transfer for the weight, making W4A16 only reaching a maximum speedup of over native FP16FP16 matrix multiplication in PyTorch. In the long run, our method lays a solid foundation and provides insightful views for the efficient deployment of quantized large language models on various domain-specific accelerators.
@@ 1 Introduction
Quantization [8] has emerged as a pivotal technique for deploying large language models (LLMs) [14, 16, 10, 1, 17] on resource-constrained hardware. Recent studies, such as SINQ [11] and QQQ [18], have achieved significant reductions in memory footprint, bandwidth requirements, and computational cost without incurring substantial losses in model accuracy. Among various quantization schemes, weight-only strategies—wherein weights are quantized to fixed-point integers while activations remain in floating-point format—have garnered growing attention due to their capability to balance efficiency and fidelity. In particular, the W4A16 configuration (4-bit weights and 16-bit activations) [3, 9] has demonstrated remarkable potential: it retains high activation precision to preserve the dynamic range, while aggressively compressing weights to minimize storage and memory access overhead. However, unlocking the full potential of W4A16 quantization necessitates specialized kernel implementations that can efficiently unpack and dequantize 4-bit weights into FP16/BF16 formats prior to matrix multiplication. General Matrix Multiplication (GEMM) routines are ill-suited for such preprocessing-intensive, irregular data layouts and often fail to fuse dequantization with computation. This deficiency results in suboptimal memory bandwidth utilization and redundant data movement—challenges that are particularly pronounced on domain-specific architectures such as Huawei’s Ascend NPU [19]. While the cube units of the Ascend NPU are highly optimized for native GEMM operations, they lack tailored support for the additional memory-bound dequantization steps required by 4-bit weight formats.
@@ 2.2 Motivation and Challenges
W4A16[3, 9] quantization (4-bit weights, 16-bit activations) offers an trade-off between model efficiency and accuracy. By storing weights in 4-bit integers, it reduces model size by approximately 4× compared to FP16 representations. However, a key implementation challenge remains: mainstream hardware—including GPUs and NPUs—lacks native support for mixed-precision arithmetic operations. In practice, 4-bit values must be packed into larger data types such as INT32 for storage and transfer, then dequant prior to computation. This dequant process introduces additional overhead and can become a notable performance bottleneck unless carefully optimized through efficient pipeline and memory access patterns. Thus, while W4A16 is promising for efficient inference, its practical speedup depends heavily on low-level kernel optimization and hardware-aware implementation.
@@ 4.1 Kernel Strategy Benchmarking
We evaluate the performance of the INT4-weight × FP16-activation matrix multiplication kernel on the Ascend 910 NPU across a range of practical matrix dimensions derived from OpenPangu, DeepSeek-R1, GLM-4.5 and LLaMA3.2 as well as different batch sizes, with a comparative analysis of two parallelization strategies: Split-K and Data Parallel. As shown in Figure 2, when is significantly larger than , the Split-K strategy outperforms data-parallel approaches, and our method achieves a speedup ranging from to . This is because Split-K can more effectively partition the computational workload across each cube core under such dimensional constraints. As the batch size (M) gradually increases, performance remains relatively stable. This occurs because the cube cores operate on matrix tiles as their fundamental processing granularity. Even when batch size is small, the input data is padded accordingly, which explains the limited acceleration benefit for small batch sizes. This behavior is observed in both Split-K and data-parallel methods.